# Electrostatics in semiconducting devices I : The Pure Electrostatics Self Consistent Approximation

### v2

@@ Abstract
preprint: APS/123-QED
@@ I Introduction
II The small parameter of the SCQE problem
@@
III The Pure Electrostatic Self-Consistent Approximation: PESCA
@@ III.1 Formulation of PESCA
The PESCA approximation amounts to taking the slope of the ILDOS to be infinite, i.e. the ILDOS is made of a horizontal branch $n(\mu<0)=0$ and a vertical branch $\mu(n>0)=0$, as shown in Fig.2c). For a bulk system it is equivalent to setting the small parameter $\kappa=0$. Keeping only the vertical branch $\mu=0$ would correspond to a good metal with a density of states high enough for the conductor to always remain an equipotential. Keeping only the horizontal branch would correspond to a fully depleted 2DEG. By keeping both branches, one allows for partially depleted devices, i.e. for a proper treatment of field-effect transistors. In particular, PESCA can predict the ”pinch-off” phase diagrams, i.e. the values of the gate voltages where the gas is depleted underneath the side gates or the middle gates in Fig.1. Solving the PESCA problem amounts to finding which part of the 2DEG is on the horizontal/vertical branch and solve the associated electrostatic problem.
@@
As argued in the preceding section, the PESCA is expected to be quantitative (within a few percent) for the calculation of thermodynamic quantities such as the density. Its level of validity for more subtle observables such as the conductance remains to be asserted. The relevance of PESCA lies in several features. (i) First, as we shall see, PESCA can be calculated very easily, at a much smaller computational cost than the full treatment of the SCQE problem. (ii) PESCA can help solving SCQE problems more accurately by treating different parts of the system at different level of description. For instance, suppose one wants to solve the SCQE problem of our split wire, c.f Fig.1, in a regime where the 2DEG is depleted below the side gates so that the wire is formed. One could treat the 2DEG inside the wire region $|x|<|x_{\rm side}|$ by solving the corresponding quantum problem. However, it is inefficient to treat the 2DEG outside of the wire region in the same way, as it would be very costly computationally. On the other hand, it is necessary to take into account that part of the 2DEG as it partially screens the effect of the gates. Treating this outside region $|x|>|x_{\rm side}|$ within PESCA would provide a precise and efficient compromise with an automatic calculation of the position of the boundary of the wire. (iii) PESCA can quantitatively (i.e. within a few percent) predict the voltages needed to deplete the 2DEG in different regions (e.g. underneath the side or middle gates). With these depletion voltages we construct what we call ”Pinch-off” phase diagrams, e.g. Fig.5. Pinch-off phase diagrams can be easily measured experimentally. They depend on the geometry of the device as well as on the charge distribution in the stack. As such, PESCA calculations may be very valuable to reconstruct this distribution and setup precise electrostatic models. Finding what is the correct distribution of e.g. ionized donors or surface charges, is indeed a prerequisite for predictive simulations of these type of devices [8]. (iv) PESCA forms the basis onto which more precise algorithms for treating Thomas-Fermi problems or SCQE problems will be built [1].
III.2 PESCA algorithm
@@
IV Application to Pinch-off phase diagrams
@@
Fig.6 shows the density at $x_{\rm wire}=-225nm$ (middle of the left subwire) as a function of the gate voltage (the same voltage is applied to all the gates). First, these curves are not linear. This is a consequence of the self-consistent nature of PESCA. Indeed for a given, fixed $\cal D/\cal N$ partitioning, the density at every point depends linearly on the gate voltage. Hence the non-linearity is a consequence of the variation of this partitioning as some regions get depleted. The different curves of Fig.6 (black, green and blue) are calculated by varying the gate voltage applied at the high temperature stage of the calculation. This is aimed at mimicking a common experimental practice known as ”bias cooling”, where a positive bias is applied during the cooling of the sample. This bias attract electrons under the gate which in turn are screened by charges elsewhere in the device. In our model, it is the surface charges that screen the electrons under the gate. At low temperature these screening charges get frozen, hence remain even if the bias voltage is removed. They contribute to depleting the 2DEG underneath the gate even at zero voltage. We find that this effect is qualitatively reproduced by our calculations. However, typically a one to one correspondence is found experimentally between the bias used during cooldown and the gate value at which pinch-off is found at low temperature [47]. In our calculations the offset in the pinch-off value is only that of half the cooldown bias voltage. This points to the presence of additional screening effects in the donor layer not taken into account here. For instance, we have supposed that the donor ionization is fully frozen at high temperature.
@@
V Adjusting the Pinch-off phase diagram to experimental data
@@
Middle & Side gates $\mathcal{D}$ $\mathcal{D}$ $\mathcal{D}$ $\mathcal{D}$ $\mathcal{D}$ $\mathcal{D}$ $\mathcal{D}$ $\mathcal{D}$
@@
VI Extension of PESCA to the integer quantum Hall effect
@@
We end this article with the first extension of the PESCA algorithm to more complex situations. As we have seen, PESCA essentially relies on the ILDOS having two different branches. However, it is straightforward to generalize the algorithm so that it can handle an ILDOS with a discrete number of branches. Such a situation arises naturally in the quantum Hall regime when the 2DEG is put under a perpendicular magnetic field. The role of electron-electron interactions is known to have drastic effects in this regime and in particular leads to the reconstruction of the edge states into compressible and incompressible stripes [3, 5, 4]. The density of states of a bulk 2DEG in the quantum Hall regime is indeed highly non-linear. It is made of delta function peaks in the (highly degenerate) Landau levels separated by insulating regions of vanishing density of states. The resulting ILDOS is a step like function as shown in Fig.10a. In Fig.10a, $\hbar\omega_{c}=2.45$ meV, with $\omega_{c}$ the cyclotron frequency, while $l_{B}=21.6$ nm is the magnetic length. To solve the self consistent problem corresponding to this ILDOS, one updates the PESCA algorithm by introducing the cells ${\cal D}_{p}$ and ${\cal N}_{p}$. The integer $p\geq 0$ indicates on which branch of the ILDOS the cell is. The ${\cal D}_{p}$ cells are Dirichlet cells with fixed potential $U=\hbar\omega_{c}(p+1/2)$ while the ${\cal N}_{p}$ cells are Neumann cells with fixed density $n=p/(2\pi l_{B}^{2})$. Step III of the PESCA algorithm also needs to be extended as follows. For the ${\cal D}_{p}$ cells,
@@
  • if $U_{i}\geq-\hbar\omega_{c}(p+1/2)$ then the cell must change it to ${\cal D}_{p-1}$.
@@
Figure 11: $\mathcal{N}_{p}/\mathcal{D}_{p}$ partitioning as a function of the iteration number for the results in Fig.12. (a) Up to iteration $12$ the PESCA approximation was used. At the latter iteration the PESCA $\mathcal{N}/\mathcal{D}$ partitioning is stable. The resulting potential and charge profile are then used as input to iteration $13$. For the latter and onwards thomas-fermi approximation is used for $B=1.42T$. The yellow and black regions correspond to $\mathcal{N}$ and $\mathcal{D}$ cells under PESCA approximation. Under thomas fermi the cells located within the first five landau levels are shown in varying degrees of green for $\mathcal{N}_{p}$ cells and blue for $\mathcal{D}_{p}$ cells. $p$ is the filling factor s.t. $n=p/(2\pi l_{b})$ and $m$ is defined s.t. $p=(1+m)/2\hbar\omega_{c}$. (b) The thomas fermi approximation for $B=1.42T$ is used for all iterations.
@@ VII Conclusion
Acknowledgements.

### v3

@@ Abstract
Copyright attribution to authors.
This work is a submission to SciPost Physics.
License information to appear upon publication.
Publication information to appear upon publication.
Received Date
Accepted Date
Published Date
@@ 1 Introduction
2 The small parameter of the SCQE problem
@@
3 The Pure Electrostatic Self-Consistent Approximation: PESCA
@@ 3.1 Formulation of PESCA
The PESCA approximation amounts to taking the slope of the ILDOS to be infinite, i.e. the ILDOS is made of a horizontal branch $n(\mu<0)=0$ and a vertical branch $\mu(n>0)=0$, as shown in Fig.2c). For a bulk system it is equivalent to setting the small parameter $\kappa=0$. Keeping only the vertical branch $\mu=0$ would correspond to a good metal with a density of states high enough for the conductor to always remain an equipotential. Keeping only the horizontal branch would correspond to a fully depleted 2DEG. By keeping both branches, one allows for partially depleted devices, i.e. for a proper treatment of field-effect transistors. In particular, PESCA can predict the "pinch-off" phase diagrams, i.e. the values of the gate voltages where the gas is depleted underneath the side gates or the middle gates in Fig.1. Solving the PESCA problem amounts to finding which part of the 2DEG is on the horizontal/vertical branch and solve the associated electrostatic problem.
@@
As argued in the preceding section, the PESCA is expected to be quantitative (within a few percent) for the calculation of thermodynamic quantities such as the density. Its level of validity for more subtle observables such as the conductance remains to be asserted. The relevance of PESCA lies in several features. (i) First, as we shall see, PESCA can be calculated very easily, at a much smaller computational cost than the full treatment of the SCQE problem. (ii) PESCA can help solving SCQE problems more accurately by treating different parts of the system at different level of description. For instance, suppose one wants to solve the SCQE problem of our split wire, c.f Fig.1, in a regime where the 2DEG is depleted below the side gates so that the wire is formed. One could treat the 2DEG inside the wire region $|x|<|x_{\rm side}|$ by solving the corresponding quantum problem. However, it is inefficient to treat the 2DEG outside of the wire region in the same way, as it would be very costly computationally. On the other hand, it is necessary to take into account that part of the 2DEG as it partially screens the effect of the gates. Treating this outside region $|x|>|x_{\rm side}|$ within PESCA would provide a precise and efficient compromise with an automatic calculation of the position of the boundary of the wire. (iii) PESCA can quantitatively (i.e. within a few percent) predict the voltages needed to deplete the 2DEG in different regions (e.g. underneath the side or middle gates). With these depletion voltages we construct what we call "Pinch-off" phase diagrams, e.g. Fig.5. Pinch-off phase diagrams can be easily measured experimentally. They depend on the geometry of the device as well as on the charge distribution in the stack. As such, PESCA calculations may be very valuable to reconstruct this distribution and setup precise electrostatic models. Finding what is the correct distribution of e.g. ionized donors or surface charges, is indeed a prerequisite for predictive simulations of these type of devices [8]. (iv) PESCA forms the basis onto which more precise algorithms for treating Thomas-Fermi problems or SCQE problems will be built [1].
3.2 PESCA algorithm
@@
4 Application to Pinch-off phase diagrams
@@
Fig.6 shows the density at $x_{\rm wire}=-225nm$ (middle of the left subwire) as a function of the gate voltage (the same voltage is applied to all the gates). First, these curves are not linear. This is a consequence of the self-consistent nature of PESCA. Indeed for a given, fixed $\cal D/\cal N$ partitioning, the density at every point depends linearly on the gate voltage. Hence the non-linearity is a consequence of the variation of this partitioning as some regions get depleted. The different curves of Fig.6 (black, green and blue) are calculated by varying the gate voltage applied at the high temperature stage of the calculation. This is aimed at mimicking a common experimental practice known as "bias cooling", where a positive bias is applied during the cooling of the sample. This bias attract electrons under the gate which in turn are screened by charges elsewhere in the device. In our model, it is the surface charges that screen the electrons under the gate. At low temperature these screening charges get frozen, hence remain even if the bias voltage is removed. They contribute to depleting the 2DEG underneath the gate even at zero voltage. We find that this effect is qualitatively reproduced by our calculations. However, typically a one to one correspondence is found experimentally between the bias used during cooldown and the gate value at which pinch-off is found at low temperature [47]. In our calculations the offset in the pinch-off value is only that of half the cooldown bias voltage. This points to the presence of additional screening effects in the donor layer not taken into account here. For instance, we have supposed that the donor ionization is fully frozen at high temperature.
@@
5 Adjusting the Pinch-off phase diagram to experimental data
@@
[…]
@@
$\dagger$ Here Gates indicate the Middle and Side gates region.
@@
6 Extension of PESCA to the integer quantum Hall effect
@@
We end this article with the first extension of the PESCA algorithm to more complex situations. As we have seen, PESCA essentially relies on the ILDOS having two different branches. However, it is straightforward to generalize the algorithm so that it can handle an ILDOS with a discrete number of branches. Such a situation arises naturally in the quantum Hall regime when the 2DEG is put under a perpendicular magnetic field. The role of electron-electron interactions is known to have drastic effects in this regime and in particular leads to the reconstruction of the edge states into compressible and incompressible stripes [3, 5, 4]. The density of states of a bulk 2DEG in the quantum Hall regime is indeed highly non-linear. It is made of delta function peaks in the (highly degenerate) Landau levels separated by insulating regions of vanishing density of states.
The resulting ILDOS is a step like function as shown in Fig.10a. In Fig.10a, $\hbar\omega_{c}=2.45$ meV, with $\omega_{c}$ the cyclotron frequency, while $l_{B}=21.6$ nm is the magnetic length. To solve the self consistent problem corresponding to this ILDOS, one updates the PESCA algorithm by introducing the cells ${\cal D}_{p}$ and ${\cal N}_{p}$. The integer $p\geq 0$ indicates on which branch of the ILDOS the cell is. The ${\cal D}_{p}$ cells are Dirichlet cells with fixed potential $U=\hbar\omega_{c}(p+1/2)$ while the ${\cal N}_{p}$ cells are Neumann cells with fixed density $n=p/(2\pi l_{B}^{2})$. Step III of the PESCA algorithm also needs to be extended as follows. For the ${\cal D}_{p}$ cells,
@@
  • if $U_{i}\geq-\hbar\omega_{c}(p+1/2)$ then the cell must change it to ${\cal D}_{p-1}$.F
@@
Figure 11: $\mathcal{N}_{p}/\mathcal{D}_{p}$ partitioning as a function of the iteration number for the results in Fig.12. Top panel: Up to iteration $12$ the PESCA approximation was used. At the latter iteration the PESCA $\mathcal{N}/\mathcal{D}$ partitioning is stable. The resulting potential and charge profile are then used as input to iteration $13$. For the latter and onwards thomas-fermi approximation is used for $B=1.42T$. The yellow and black regions correspond to $\mathcal{N}$ and $\mathcal{D}$ cells under PESCA approximation. Under thomas fermi the cells located within the first five landau levels are shown in varying degrees of green for $\mathcal{N}_{p}$ cells and blue for $\mathcal{D}_{p}$ cells. $p$ is the filling factor s.t. $n=p/(2\pi l_{b})$ and $m$ is defined s.t. $p=(1+m)/2\hbar\omega_{c}$. Bottom panel : The thomas fermi approximation for $B=1.42T$ is used for all iterations.
@@ 7 Conclusion
Acknowledgements
We express our gratitude to Christoph Groth for insightful programming advice. We thank Christopher Bäuerle, Frederic Pierre and Patrice Roche for helpful discussions.
Funding information